# Linear Complementary Pairs of Quasi-Cyclic and Quasi-Twisted Codes

### v1

@@ 4 LCPs of one-generator quasi-cyclic codes of index 2
We show that $(i),(ii),(iii),(iv)$ are equivalent to $(A),(B),(C)$. First, assume that $(i)$, $(ii)$, $(iii)$, and $(iv)$ hold. Then, by Lemma 4.2 $(C)$ holds. Also, $(ii),(iii)$ are the same as $(A)$ and $(B)$.
Conversely, suppose $(A),(B),(C)$ hold. Then $(ii)$, $(iii)$ and $(iv)$ hold trivially. If $\gcd(g_{11}(x),f_{11}(x))\neq 1$, then $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))\neq 1$. Hence $(C)$ implies $(i)$. This completes the proof. ∎
In [5], the authors called a code $C$ maximal if $\dim(C)=m$. Recall that if $C$ is a one-generator quasi-cyclic code of length $2m$ and index $2$, generated by $(g_{11}(x),g_{12}(x))$ with $g_{11}(x)\mid(x^{m}-1)$, then $\dim(C)=m-\deg\ g(x)$, where $g(x)=\gcd(g_{11}(x),g_{12}(x))$.
Corollary 4.4.
[5, Proposition 3.6] Let $C=\langle(g_{11}(x),g_{12}(x))\rangle$ and $D=\langle(f_{11}(x),f_{12}(x))\rangle$ be maximal one-generator quasi-cyclic codes of index $2$, where $g_{11}(x),f_{11}(x)\mid(x^{m}-1)$. Then $(C,D)$ is an LCP of codes if and only if $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$.
Let $R=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$. Double circulant (DC) codes are one-generator quasi-cyclic codes of index $2$, generated by $\langle(1,a(x))\rangle\in R^{2}$. In the next result, we show that Theorem 4.3 generalizes [9, Proposition 3.2].
@@ 5 LCPs of quasi-twisted codes of index 2
Let $C$ be a $\lambda$-quasi-twisted code of length $2m$ and index $2$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
@@
Moreover, in this case $\dim C=2m-\deg g_{11}(x)-\deg g_{22}(x)$.
@@
In [16] quasi-cyclic codes (i.e., quasi-twisted codes with $\lambda=1$) were studied using Gröbner bases of modules. We note that the results of [16] are valid for any $\lambda$ (all proofs are valid if we take $x^{m}-\lambda$ instead of $x^{m}-1$). Hence, by [16] we can assume that $C$ is generated by two elements $g_{1}=\big{(}g_{11}(x),g_{12}(x)\big{)}$ and $g_{2}=\big{(}0,g_{22}(x)\big{)}$, satisfying the conditions $g_{11}(x)\mid(x^{m}-\lambda)$, $g_{22}(x)\mid(x^{m}-\lambda)$, $\deg g_{12}(x)<\deg g_{22}(x)$, and $\dim C=2m-\deg g_{11}(x)-\deg g_{22}(x)$. The existence of such generators is equivalent [16] to the existence of a $2\times 2$ polynomial matrix $(a_{ij})$ such that
@@
Let $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. Then, by Remark 5.2, $g(x)\mid g_{12}(x)$ and $f(x)\mid f_{12}(x)$. Let $g_{11}(x)=g(x)g_{11}^{\prime}(x),g_{22}=g(x)g_{22}^{\prime}(x)$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x),f_{22}=f(x)f_{22}^{\prime}(x)$. Next, we provide another characterization of two-generator and one-generator LCPs of quasi-twisted codes of index $2$.
@@
[…]
@@
Let $C$ and $D$ be one-generator $\lambda$-quasi-twisted codes generated by $(g_{11}(x),g_{12}(x))$, and $(f_{11}(x),f_{12}(x))$ respectively, where $g_{11}(x),f_{11}(x)\mid(x^{m}-\lambda)$. Then $(C,D)$ is an LCP of codes if and only if
  1. (A)
    $\gcd(g_{11}(x),g_{12}(x))=1$.
  2. (B)
    $\gcd(f_{11}(x),f_{12}(x))=1$.
  3. (C)
    $\gcd(x^{m}-\lambda,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x)=1$.
@@ 6 Examples of LCPs of codes with good security parameters
In this section, we present several examples of LCPs of codes with security parameters. The parameters of a linear code are denoted by $[n,k,d]$, where $n$ is the length, $k$ is the dimension, and $d$ is the minimum Hamming distance of the code. We denote the security parameter of an LCP of codes $(C,D)$ by $d_{LCP}=\min\{d(C),d(D^{\perp})\}$ for given values of $n,k$, and $d_{BKLC}$ denotes the minimum distance of the best known linear code for given $n,k$. All computations are done in MAGMA [7]. For the values of $d_{BKLC}$, we refer to Grassl’s code table [11].
@@
$\begin{split}g_{11}(x)&=x^{2}+3x+2,\\ g_{12}(x)&=x^{5}+4x^{4}+4x^{3}+x^{2}+4x+4,\\ g_{22}(x)&=x^{6}+x^{5}+3x^{4}+x^{2}+x+3\end{split}$
@@
$\begin{split}f_{11}(x)&=x+3,\\ f_{12}(x)&=x^{3}+2x^{2}+x+2,\\ f_{22}(x)&=x^{7}+3x^{6}+4x^{5}+2x^{4}+x^{3}+3x^{2}+4x+2.\end{split}$
Then $(C,D)$ is an LCP of quasi-cyclic codes with $n=16$, $k=8$ and $d_{LCP}=6$. Note that for $n=16$, $k=8$, we have $d_{BKLC}=7$.
@@
$\begin{split}g_{11}(x)&=x+1,\\ g_{12}(x)&=x^{2}+x+1,\\ g_{22}(x)&=x^{3}+2x^{2}+x+2\end{split}$
@@
$\begin{split}f_{11}(x)&=x+2,\\ f_{12}(x)&=x^{2}+x+2,\\ f_{22}(x)&=x^{3}+x^{2}+x+1.\end{split}$
Then $(C,D)$ is an LCP of ternary quasi-cyclic codes with $n=8$, $k=4$ and $d_{LCP}=4$. This security parameter is equal to the minimum distance $d_{BKLC}$ of the best known linear code.
@@
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{2}+2x+2,\\ g_{22}(x)&=x^{4}+x^{3}+x^{2}+x+1\end{split}$
@@
Then $(C,D)$ is an LCP of ternary quasi-cyclic codes with $n=10$, $k=5$ and $d_{LCP}=5=d_{BKLC}$. This security parameter improves the security parameter obtained in [5, Table 3] with $n=10$ $k=5$.
In Table 1, we list some examples of LCPs of quasi-cyclic codes $(C,D)$ over $\mathbb{F}_{4}=\mathbb{F}_{2}[w]$, where the code $C$ is generated by $(g_{11}(x),g_{12}(x)),(0,g_{22}(x))$ and the code $D$ is generated by $(f_{11}(x),f_{12}(x)),(0,f_{22}(x))$. We compare the security parameter $d_{LCP}$ with $d_{BKLC}$.
$n$ $k$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ $f_{11}(x)$ $f_{12}(x)$ $f_{22}(x)$ $d_{LCP}$ $d_{BKLC}$
$6$ $3$ $x+w$ $x+1+w^{2}$ $x^{2}+wx$ $+w^{2}$ $1$ $wx^{3}+wx^{2}$ $+w^{2}x+1$ $x^{3}-1$ $3$ $4$
$10$ $5$ $x+1$ $x^{3}+wx^{2}$ $+wx$ $x^{4}+x^{3}$ $+x^{2}+x$ $+1$ $1$ $x^{4}+wx^{3}$ $+x+w$ $x^{5}-1$ $5$ $5$
$14$ $9$ $x+1$ $x^{3}+w^{2}x$ $+w$ $x^{4}+x^{3}$ $+x^{2}+1$ $x^{3}+x^{2}$ $+1$ $x^{3}+x^{2}$ $+1$ $x^{6}+x^{5}$ $+x^{4}+x^{3}$ $+x^{2}+x+1$ $4$ $4$
Table 1: LCPs of quasi-cyclic codes over $\mathbb{F}_{4}$
@@
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{6}+4x^{5}+2x^{3}+3x^{2}+x+3,\\ g_{22}(x)&=x^{10}+3x^{9}+4x^{8}+2x^{7}+x^{6}+3x^{5}+4x^{4}+2x^{3}+x^{2}+3x+4\end{split}$
@@
Then $(C,D)$ is an LCP of $\lambda$-quasi-twisted quinary codes with $n=22$, $k=11$ and $d_{LCP}=8$. This security parameter $d_{LCP}$ is the same as the minimum distance of the best known linear codes $d_{BKLC}=8$.
@@
Let $q=9,$ $m=10$ and $\lambda=w$, where $\mathbb{F}_{9}=\mathbb{F}_{3}[w]$ and $w^{2}+2w+2=0$. Suppose $C$ is a $\lambda$-quasi-twisted code generated by
$\begin{split}g_{11}(x)&=x^{2}+2x+w,\\ g_{12}(x)&=x^{5}+w^{2}x^{4}+w^{5}x^{3}+w^{6}x^{2}+w^{2}x+w^{3},\\ g_{22}(x)&=x^{8}+x^{7}+w^{3}x^{6}+w^{2}x^{5}+w^{7}x^{4}+w^{3}x^{3}+w^{5}x^{2}+w^{3}x+2\end{split}$
and $D$ is an another $\lambda$-quasi-twisted code generated by
$\begin{split}f_{11}(x)&=x^{2}+w^{7}x+w,\\ f_{12}(x)&=x^{7}+w^{7}x^{6}+wx^{5}+w^{6}x^{4}+x^{3}+w^{7}x^{2}+w^{3}x,\\ f_{22}(x)&=x^{8}+w^{3}x^{7}+w^{7}x^{6}+w^{7}x^{5}+wx^{4}+x^{3}+wx^{2}+w^{6}x+2.\end{split}$
Then $(C,D)$ is an LCP of $\lambda$-quasi-twisted codes over $\mathbb{F}_{9}$ with $n=20$, $k=10$ and $d_{LCP}=8$. The minimum distance of the best known linear codes is $d_{BKLC}=10$.
@@
Let $q=9,$ $m=10$ and $\lambda=w$, where $\mathbb{F}_{9}=\mathbb{F}_{3}[w]$ and $w^{2}+2w+2=0$. Suppose $C$ is a $\lambda$-quasi-twisted code generated by
$\begin{split}g_{11}(x)&=1,\\ g_{12}(x)&=x^{5}+x^{4}+wx^{3}+w^{5}x^{2}+w^{2},\\ g_{22}(x)&=x^{6}+w^{3}x^{5}+2x^{4}+wx^{3}+w^{5}x^{2}+w^{5}x+w^{3}\end{split}$
and $D$ is an another $\lambda$-quasi-twisted code generated by
$\begin{split}f_{11}(x)&=x^{4}+w^{7}x^{3}+w^{5}x^{2}+x+w^{2},\\ f_{12}(x)&=w^{6}x^{8}+w^{3}x^{7}+w^{7}x^{6}+w^{6}x^{5}+w^{7}x^{4}+w^{3}x^{3}+x^{2}+w^{5}x,\\ f_{22}(x)&=x^{10}+w^{5}.\end{split}$
Then $(C,D)$ is an LCP of $\lambda$-quasi-twisted codes over $\mathbb{F}_{9}$, where $C$ is a $[20,14]$ code and $d_{LCP}=5$. The minimum distance of the best known linear codes with $n=20$ and $k=14$ is $d_{BKLC}=6$.

### v2

@@ 4 LCPs of one-generator quasi-cyclic codes of index 2
We show that $(i),(ii),(iii)$ are equivalent to
$\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1.$ (4.2)
First, assume that $(i)$, $(ii)$, and $(iii)$ hold. Then, by Lemma 4.2, $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$ holds.
Conversely, if $\gcd(g_{11}(x),f_{11}(x))\neq 1$ or $g(x)\neq 1$ or $f(x)\neq 1$, then $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))\neq 1$. Hence, Eq. 4.2 implies $(i)$, and $(ii)$. Moreover, by Lemma 4.2, Eq. 4.2 along with $(i)$ and $(ii)$, implies $(iii)$. This completes the proof. ∎
Remark 4.1.
The necessary and sufficient condition in Theorem 4.3, $\gcd(x^{m}-1,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$, can be viewed as an analogue of the vector space case. Let $C$ and $D$ be one-generator quasi-cyclic codes generated by $(g_{11}(x),g_{12}(x))$ and $(f_{11}(x),f_{22}(x))$, respectively. Then $(C,D)$ is an LCP of codes if and only if the $2\times 2$ matrix
$\begin{bmatrix}g_{11}(x)&g_{12}(x)\\ f_{11}(x)&f_{12}(x)\end{bmatrix}$
is invertible.
Let $R=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$. Double circulant (DC) codes are one-generator quasi-cyclic codes of index $2$, generated by $\langle(1,a(x))\rangle\in R^{2}$. In the next result, we show that Theorem 4.3 generalizes [10, Proposition 3.2].
@@ 5 LCPs of quasi-twisted codes of index 2
Let $C$ be a $\lambda$-quasi-twisted code of length $2m$ and index $2.$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
@@
Moreover, in this case $\dim(C)=2m-\deg g_{11}(x)-\deg g_{22}(x)$.
@@
In [20] quasi-cyclic codes (i.e., quasi-twisted codes with $\lambda=1$) were studied using Gröbner bases of modules. We note that the results of [20] are valid for any $\lambda$ (all proofs are valid if we take $x^{m}-\lambda$ instead of $x^{m}-1$). Hence, by [20] we can assume that $C$ is generated by two elements $g_{1}=\big(g_{11}(x),g_{12}(x)\big)$ and $g_{2}=\big(0,g_{22}(x)\big)$, satisfying the conditions $g_{11}(x)\mid(x^{m}-\lambda)$, $g_{22}(x)\mid(x^{m}-\lambda)$, $\deg g_{12}(x)<\deg g_{22}(x)$, and $\dim(C)=2m-\deg g_{11}(x)-\deg g_{22}(x)$. The existence of such generators is equivalent [20] to the existence of a $2\times 2$ polynomial matrix $(a_{ij})$ such that
@@
Let $g(x)=\gcd(g_{11}(x),g_{22}(x))$ and $f(x)=\gcd(f_{11}(x),f_{22}(x))$. Then, by Remark 5.1, $g(x)\mid g_{12}(x)$ and $f(x)\mid f_{12}(x)$. Let $g_{11}(x)=g(x)g_{11}^{\prime}(x),g_{22}(x)=g(x)g_{22}^{\prime}(x)$, and $f_{11}(x)=f(x)f_{11}^{\prime}(x),f_{22}(x)=f(x)f_{22}^{\prime}(x)$. Next, we provide another characterization of two-generator and one-generator LCPs of quasi-twisted codes of index $2$.
@@
Let $C$ and $D$ be $\lambda$-quasi-twisted codes of length $2m$ and index $2$ generated by $(g_{11}(x),g_{12}(x))$, $(0,g_{22}(x))$ and $(f_{11}(x),f_{12}(x))$, $(0,f_{22}(x))$, respectively, satisfying Condition ($\ast\ast$). Then $(C,D)$ is an LCP of quasi-twisted codes if and only if
@@
Let $C$ and $D$ be one-generator $\lambda$-quasi-twisted codes generated by $(g_{11}(x),g_{12}(x))$, and $(f_{11}(x),f_{12}(x))$ respectively, where $g_{11}(x),f_{11}(x)\mid(x^{m}-\lambda)$. Then $(C,D)$ is an LCP of codes if and only if $\gcd(x^{m}-\lambda,g_{11}(x)f_{12}(x)-g_{12}(x)f_{11}(x))=1$.
@@ 6 Examples of LCPs of codes with good security parameters
The parameters of a linear code are denoted by $[n,k,d]$, where $n$ is the length, $k$ is the dimension, and $d$ is the minimum Hamming distance of the code. For given $n,k$, we denote the security parameter of an LCP of codes $(C,D)$ by $d_{LCP}(C,D)=\min\{d(C),d(D^{\perp})\}$, the optimal distance of linear codes by $d_{opt}$, that is, any $n,k,>d_{opt}]$ code does not exist, and $d_{BKLC}$ denotes the minimum distance of the best known linear code. For the values of $d_{opt}$ and $d_{BKLC}$, we refer to Grassl’s code table [14]. In [11], Carlet et al. proved that for $q>2$, there exists an LCP of linear codes with $d_{LCP}(C,D)=d_{opt}$, and for $q=2$, LCPs of linear codes exist with $d_{LCP}(C,D)\geq d_{opt}-1$. In this section, we illustrate that the algebraic characterizations in the preceding sections can be utilized to construct LCPs of quasi-cyclic and quasi-twisted codes with $d_{LCP}(C,D)=d_{opt}$ ($q>2$), and $d_{LCP}(C,D)=d_{opt}-1$ or $d_{LCP}(C,D)=d_{opt}\ (q=2)$. All computations are done in MAGMA [7].
@@
$\begin{split}g_{11}(x)&=(x+2),\\ g_{12}(x)&=(x+2)((x^{3}+x^{2}+1)),\\ g_{22}(x)&=(x+1)(x+2)(x+4)(x^{2}+2)(x^{2}+3),\end{split}$
@@
$\begin{split}f_{11}(x)&=(x+3),\\ f_{12}(x)&=(x+3)(4x^{3}+x^{2}+1),\\ f_{22}(x)&=(x+1)(x+3)(x+4)(x^{2}+2)(x^{2}+3).\end{split}$
Then, observe that
  1. 1.
    $\gcd(g_{11}(x),f_{11}(x))=1$.
  2. 2.
    $g_{11}(x)g_{22}(x)f_{11}(x)f_{22}(x)=(x^{8}-1)^{2}$.
  3. 3.
Thus, by Theorem 3.2, $(C,D)$ is an LCP of quasi-cyclic codes with $n=16$, $k=8$, and $d_{LCP}(C,D)=7=d_{opt}$. Moreover, a cyclic code with these parameters does not exist.
@@
$\begin{split}g_{11}(x)&=x+1,\\ g_{12}(x)&=x^{2}+x+1,\\ g_{22}(x)&=(x+2)(x^{2}+1),\end{split}$
@@
$\begin{split}f_{11}(x)&=x+2,\\ f_{12}(x)&=x^{2}+x+2,\\ f_{22}(x)&=(x+1)(x^{2}+1).\end{split}$
Then, by Theorem 3.2, $(C,D)$ is an LCP of ternary quasi-cyclic codes with $n=8$, $k=4$, and $d_{LCP}(C,D)=4=d_{opt}$.
@@
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{2}+2x+2,\\ g_{22}(x)&=x^{4}+x^{3}+x^{2}+x+1,\end{split}$
@@
Then $(C,D)$ is an LCP of ternary quasi-cyclic codes with $n=10$, $k=5$ and $d_{LCP}(C,D)=5=d_{opt}$. This security parameter improves the security parameter obtained in [5, Table 3] with $n=10$ and $k=5$.
In Table 1, we list some examples of LCPs of quasi-cyclic codes $(C,D)$ over $\mathbb{F}_{4}=\mathbb{F}_{2}[w]$, where the code $C$ is generated by $(g_{11}(x),g_{12}(x)),(0,g_{22}(x))$ and the code $D$ is generated by $(f_{11}(x),f_{12}(x)),(0,f_{22}(x))$. We compare the security parameter $d_{LCP}(C,D)$ with $d_{BKLC}$.
$n$ $k$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ $f_{11}(x)$ $f_{12}(x)$ $f_{22}(x)$ $d_{LCP}$ $d_{opt}$
$10$ 55 $x+1$ $x^{3}+wx^{2}$ +wx+wx $x^{4}+x^{3}$ $x^{2}+x +$1+1 11 $x^{4}+wx^{3}$ +x+w+x+w $x^{5}-1$ 55 55
$14$ 99 $x+1$ $x^{3}+w^{2}x$ +w+w $x^{4}+x^{3}$ $x^{2}+1$ $x^{3}+x^{2}$ +1+1 $x^{3}+x^{2}$ +1+1 $x^{6}+x^{5}$ $x^{4}+x^{3} +$$x^{2}+x+1$ 44 44
Table 1: LCPs of quasi-cyclic codes over $\mathbb{F}_{4}$.
@@
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{6}+4x^{5}+2x^{3}+3x^{2}+x+3,\\ g_{22}(x)&=(x^{5}+x^{4}+x^{3}+2x^{2}+x+2)(x^{5}+2x^{4}+x^{3}+2x^{2}+3x+2),\end{split}$
@@
Then, by Theorem 5.3, $(C,D)$ is an LCP of $\lambda$-quasi-twisted quinary codes with $n=22$, $k=11$, and $d_{LCP}(C,D)=8=d_{BKLC}$.
In [8], the author showed that $d_{LCD}(C)\leq 7$ for $n=30$ and $k=15$ over $\mathbb{F}_{2}$, where $d_{LCD}(C)$ is the minimum distance of a binary LCD code $C$. In the following example, we show that there is an LCP of codes with $d_{LCP}(C,D)=7$ for the same $n$ and $k$.
@@
Let $q=2,$ $m=15$. Suppose $C$ is a quasi-cyclic code generated by
$\begin{split}g_{11}(x)&=x^{2}+x+1,\\ g_{12}(x)&=x^{12}+x^{11}+x^{9}+x^{8}+x^{7}+x^{5}+x^{4}+x^{2},\\ g_{22}(x)&=(x^{4}+x^{3}+1)(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x+1)(x+1),\end{split}$
and $D$ is an another quasi-cyclic code generated by
$\begin{split}f_{11}(x)&=x+1,\\ f_{12}(x)&=x^{12}+x^{10}+x^{8}+x^{7}+x^{6}+x+1,\\ f_{22}(x)&=(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x+1)(x^{4}+x^{3}+1)(x^{2}+x+1).\end{split}$
Then, by Theorem 3.2, $(C,D)$ is a binary LCP of quasi-cyclic codes with $d_{LCP}(C,D)=7=d_{opt}-1$, where $C$ is a binary $[30,15]$ code.
@@
Let $q=2,$ $m=15$. Suppose $C$ is a quasi-cyclic code generated by
$\begin{split}g_{11}(x)&=x+1,\\ g_{12}(x)&=(x+1)(x^{6}+x^{5}+x^{2}+1),\\ g_{22}(x)&=x^{m}-1,\end{split}$
and $D$ is another quasi-cyclic code generated by
$\begin{split}f_{11}(x)&=1,\\ f_{12}(x)&=x^{7}+x^{5}+x^{4}+x^{3}+1,\\ f_{22}(x)&=(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x+1)(x^{4}+x^{3}+1)(x^{2}+x+1).\end{split}$
Then, by Theorem 3.2, $(C,D)$ is a binary LCP of quasi-cyclic codes with $d_{LCP}(C,D)=8=d_{opt}$, where $C$ is a binary $[30,14]$ code. This value of $d_{LCP}(C,D)$ is better than the lower bound $d_{opt}-1$.
Example 6.7.
Let $q=2,$ $m=21$. Suppose $C$ is a quasi-cyclic code generated by
$\begin{split}g_{11}(x)&=(x+1)(x^{3}+x^{2}+1)(x^{6}+x^{4}+x^{2}+x+1),\\ g_{12}(x)&=(x+1)^{3}(x^{3}+x^{2}+1)^{2}(x^{5}+x^{2}+1)(x^{6}+x^{4}+x^{2}+x+1),\\ g_{22}(x)&=x^{m}-1,\end{split}$
and $D$ is another quasi-cyclic code generated by
$\begin{split}f_{11}(x)&=1,\\ f_{12}(x)&=x^{9}+x^{6}+x^{3}+x+1,\\ f_{22}(x)&=(x^{2}+x+1)(x^{3}+x+1)(x^{6}+x^{5}+x^{4}+x^{2}+1).\end{split}$
Then $(C,D)$ is a binary LCP of quasi-cyclic codes with $d_{LCP}(C,D)=16=d_{opt}$, where $C$ is a binary $[42,11,16]$ code. This value of $d_{LCP}(C,D)$ is better than the lower bound $d_{opt}-1$.